# Role-Selection Game in Block Production under Proposer-Builder Separation

### v1

@@ III-C Empirical Game-Theoretic Analysis
IV Discussion
@@
Using game theory to study MEV is crucial for understanding participant behavior and mitigating negative externalities in blockchain systems, as highlighted by the formalization of MEV games and the comparison of transaction ordering mechanisms in related work [19]. Due to the involvement of agents employing two types of strategies and the intricacies of the block-building process, the proposed game-theoretic problem is difficult to solve for Nash equilibrium using classical methods. We adopt agent-based simulation methods and genetic algorithms to model the co-evolution of agents, addressing the challenge highlighted in [20] regarding agent-based modeling in DeFi systems.
@@
This paper primarily focuses on the strategy selection and value allocation of profit-seeking actors involved in the block-building process. We assume that builders will use the blocks they construct to participate in an equivalent of a second-price auction, without considering any strategic behavior by builders in the block-building auction. However, a more interesting scenario is that each agent participates as a block builder in the block building auction, while they can also choose to share their bundles to reduce risk. This leads to complex bidding strategies in the block building auction. Recently, Wu et al. [21, 22] construct an agent-based model to perform empirical game-theoretic analysis on the bidding strategies of builders in the block building auction associated with the PBS system. Combining models of these two distinct stages of the MEV supply chain would induce a more complex game-theoretic problem but could offer deeper insights into understanding PBS systems.
V Related Works
V-A Transaction Fee Mechanisms
Certain researchers employ mechanism design theory to evaluate the feasibility of establishing a dominant-strategy incentive-compatible transaction fee mechanism [23, 24, 25, 26, 27]. However, the impossibility results derived from these theoretical studies indicate that the desired attributes of an ideal transaction fee mechanism cannot be achieved simultaneously. For instance, Bahrani et al. illustrate that no fee mechanism can attain incentive compatibility if block producers strive to actively extract MEV, as commonly observed in practice [26]. The analysis in [25], focused on a transaction fee mechanism (TFM) problem for a model featuring searchers and block producers (builder-proposer integration in PBS), considers a more detailed block production process, particularly a decentralized version of PBS. Unlike the literature grounded in algorithmic game theory or mechanism design, this paper primarily focuses on the strategies of economic agents within the MEV supply chain under the more practical framework of non-revealing mechanisms, specifically the first-price auction.
V-B MEV Supply Chain
Bahrani et al. [5] conduct an assessment of the effectiveness of PBS, with results quantifying, as a function of the competitiveness of the builder ecosystem, the extent to which PBS mitigates reward heterogeneity among different proposers. In recent literature, several studies have investigated the PBS landscape, underscoring the mounting tendency towards centralization of block building within the framework of PBS. Gupta et al. [28] empirically demonstrate that builders proficient in capturing CEX/DEX arbitrage (e.g., HFT firms) can construct blocks with significant top-of-block value. Further, the authors prove that builders with superior top-of-block capabilities are predisposed to dominate order flow auctions (OFAs) and subsequently exploit the private order flow acquired in these OFAs to dominate the PBS auction. Capponi et al. [6] propose a three-stage game-theoretical model comprising heterogeneous block builders and a singular order flow provider. The subgame perfect equilibrium illustrates a positive feedback loop similar to that described in [28]: an advantage in block building auctions translates into an advantage in acquiring order flow, which, in turn, increases the likelihood of winning future block building auctions, thereby further consolidating control over order flow. Öz et al. [8] reveal that builders’ profitability is correlated with access to exclusive order flow, which is subsequently linked to their market share, thereby naturally underscoring a ”chicken-and-egg” dilemma. Yang et al. [9] propose two metrics to quantify the competitiveness and efficiency of the MEV-boost auction, highlighting issues related to current PBS, including entry barriers for builders and trust crises faced by searchers. Our work seeks to explore the landscape of MEV actors within a decentralized version of PBS, extending beyond the issue of trust. Additionally, Mamageishvili et al. [17] examine how competition among searchers influences the allocation of value between a single validator and the searchers.
@@ VI Conclusion
This study introduces a co-evolutionary framework to analyze the strategic behavior of profit-seeking participants in the block-building process within the Proposer-Builder Separation (PBS) system. By employing genetic algorithms, we simulate the co-evolution of agents’ strategies, enabling us to observe their behavioral dynamics under varying conditions, particularly changes in the probabilities of conflicts between bundles. Through agent-based simulations, we capture the evolutionary outcomes of the system and apply empirical game-theoretic methods, specifically the $\alpha$-Rank algorithm, to compute the selection frequencies of agents for two key meta-strategies: block building and bundle sharing.
Appendix A One-sided Market
To better understand the block building model, we begin by conducting static analysis on an example of a one-sided market. Consider a model with $|\mathcal{A}|=3$, and due to the intense conflict $\varphi_{1,2}=\varphi_{2,1}=-1$ between the two agents $a_{1}$ and $a_{2}$ acting as builders, they are incapable of sharing bundles with each other. We assume that the strategy of $a_{3}$ is fixed at bundle sharing, i.e., $a_{3}$ functions as a searcher. Value of each bundles is drawn from the distribution function $v_{i}\sim F_{i}$ and agents are unaware of each other’s private valuation of bundles, but they can know each other’s value distribution.
$\pi_{1}(v_{2},v_{3})=\begin{cases}(1-\alpha_{13})(v_{1}-v_{2}+(\beta_{31}-\beta_{32})v_{3}),&v_{1}+\beta_{31}v_{3}\geq v_{2}+\beta_{32}v_{3},\\ 0,&v_{1}+\beta_{31}v_{3}<v_{2}+\beta_{32}v_{3}.\end{cases}$
$\pi_{2}(v_{1},v_{3})=\begin{cases}(1-\alpha_{23})(v_{2}-v_{1}+(\beta_{32}-\beta_{31})v_{3}),&v_{2}+\beta_{32}v_{3}\geq v_{1}+\beta_{31}v_{3},\\ 0,&v_{2}+\beta_{32}v_{3}<v_{1}+\beta_{31}v_{3}.\end{cases}$
$\pi_{3}(v_{1},v_{2})=\begin{cases}(1-\beta_{31})v_{3}+\alpha_{13}(v_{1}-v_{2}+(\beta_{31}-\beta_{32})v_{3}),&v_{1}+\beta_{31}v_{3}\geq v_{2}+\beta_{32}v_{3},\\ (1-\beta_{32})v_{3}+\alpha_{23}(v_{2}-v_{1}+(\beta_{32}-\beta_{31})v_{3}),&v_{1}+\beta_{31}v_{3}<v_{2}+\beta_{32}v_{3}.\end{cases}$
Denote $\mathrm{\Delta}v=v_{1}-v_{2}$ and $\mathrm{\Delta}\beta=\beta_{31}-\beta_{32}$, we have
$\pi_{3}(\mathrm{\Delta}v)=\begin{cases}v_{3}-(\beta_{32}+\mathrm{\Delta}\beta)v_{3}+\alpha_{13}(\mathrm{\Delta}v+\mathrm{\Delta}\beta v_{3}),&\mathrm{\Delta}v\geq-\mathrm{\Delta}\beta v_{3},\\ v_{3}-\beta_{32}v_{3}+\alpha_{23}(-\mathrm{\Delta}v-\mathrm{\Delta}\beta v_{3}),&\mathrm{\Delta}v<-\mathrm{\Delta}\beta v_{3}.\end{cases}$
$f_{\mathrm{\Delta}v}(x)=\begin{cases}\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}e^{-\lambda_{1}x},&x\geq 0,\\ \frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}e^{\lambda_{2}x},&x<0\end{cases}$
$\mathbb{E}[\pi_{3}]=$ $\int_{-\mathrm{\Delta}\beta v_{3}}^{+\infty}[v_{3}-(\beta_{32}+\mathrm{\Delta}\beta)v_{3}+\alpha_{13}(x+\mathrm{\Delta}\beta v_{3})]f_{\mathrm{\Delta}v}(x)\mathrm{d}x$
$\leq$ $\int_{-\mathrm{\Delta}\beta v_{3}}^{+\infty}[v_{3}-\mathrm{\Delta}\beta v_{3}+\alpha_{13}(x+\mathrm{\Delta}\beta v_{3})]f_{\mathrm{\Delta}v}(x)\mathrm{d}x$
$+\int_{-\infty}^{-\mathrm{\Delta}\beta v_{3}}[v_{3}+\alpha_{23}(-x-\mathrm{\Delta}\beta v_{3})]f_{\mathrm{\Delta}v}(x)\mathrm{d}x$
$\mathbb{E}[\pi_{3}]=$ $\int_{0}^{+\infty}[v_{3}-\mathrm{\Delta}\beta v_{3}+\alpha_{13}(x+\mathrm{\Delta}\beta v_{3})]\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}e^{-\lambda_{1}x}\mathrm{d}x$
$+\int_{-\infty}^{-\mathrm{\Delta}\beta v_{3}}[v_{3}+\alpha_{23}(-x-\mathrm{\Delta}\beta v_{3})]\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}e^{\lambda_{2}x}\mathrm{d}x,$
it yields that
$\mathbb{E}[\pi_{3}]=$ $\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}\left(\frac{v_{3}-\mathrm{\Delta}\beta v_{3}+\alpha_{13}\mathrm{\Delta}\beta v_{3}}{\lambda_{1}}+\frac{\alpha_{13}}{\lambda_{1}^{2}}\right)$
$+\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}\left(\frac{(v_{3}-\mathrm{\Delta}\beta v_{3}+\alpha_{13}\mathrm{\Delta}\beta v_{3})(1-e^{-\lambda_{2}\mathrm{\Delta}\beta v_{3}})}{\lambda_{2}}\right)$
$+\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}\left(\frac{\alpha_{13}\mathrm{\Delta}\beta v_{3}e^{-\lambda_{2}\mathrm{\Delta}\beta v_{3}}}{\lambda_{2}}-\frac{\alpha_{13}(1-e^{-\lambda_{2}\mathrm{\Delta}\beta v_{3}})}{\lambda_{2}^{2}}\right)$
$+\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}\left(\frac{(v_{3}-\alpha_{23}\mathrm{\Delta}\beta v_{3})e^{-\lambda_{2}\mathrm{\Delta}\beta v_{3}}}{\lambda_{2}}+\frac{\alpha_{23}\mathrm{\Delta}\beta v_{3}e^{-\lambda_{2}\mathrm{\Delta}\beta v_{3}}}{\lambda_{2}}+\frac{\alpha_{23}e^{-\lambda_{2}\mathrm{\Delta}\beta v_{3}}}{\lambda_{2}^{2}}\right).$
Computing the derivative of $\mathbb{E}[\pi_{3}]$ with respect to $\mathrm{\Delta}\beta$, we can obtain
$\frac{\mathrm{d}\mathbb{E}[\pi_{3}]}{\mathrm{d}\mathrm{\Delta\beta}}$
$=$ $\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}+\lambda_{2}}\left(\frac{v_{3}(\alpha_{13}-1)}{\lambda_{1}}+\frac{v_{3}(\alpha_{13}-1)}{\lambda_{2}}\left(1-e^{-\lambda_{2}\mathrm{\Delta}\beta v_{3}}\right)-v_{3}^{2}\mathrm{\Delta}\beta e^{-\lambda_{2}\mathrm{\Delta}\beta v_{3}}-\frac{\alpha_{23}v_{3}e^{-\lambda_{2}\Delta\beta v_{3}}}{\lambda_{2}}\right)$
$<$ $0$
This implies that the searcher should adopt a $\mathrm{\Delta}\beta=0$ strategy, meaning uniform bidding to all builders.

### v2

@@ IV-C Empirical Game-Theoretic Analysis
Fig. 5 clearly demonstrates that an increase in the probability of conflicts leads agents to prefer the block building strategy, which entails building blocks to engage in competition within the block building auction, instead of resorting to the bundle sharing strategy where they share their bundles with builders to evade fierce competition. This result aligns well with intuitive reasoning. In particular, we can identify the critical probability of conflicts at which agents’ strategies transition, which is approximately $0.2$.
Our results show that when the probability of bundle conflicts is low, searchers capture the majority of value within the bundles, and builders must actively participate in block-building auctions to extract additional profits. As the likelihood of conflicts increases, however, competition among searchers intensifies, prompting higher bid ratios to ensure inclusion. This heightened competition shifts more value from searchers to builders, as increased bids enhance the value provided to builders. In response, builders strategically adjust their rebate ratios to incentivize searchers to differentiate their bids, thereby improving the builders’ own chances in subsequent block-building auctions. However, our analysis also reveals that this negative feedback—where builders influence searcher bidding through rebate adjustments—is secondary to the direct effect of increased searcher competition. Ultimately, as competition escalates, remaining a searcher becomes less profitable, making the builder role increasingly attractive. These findings underscore that bundle conflicts serve as a key mechanism driving dynamic role-switching behavior within the system.
V Discussion
@@
Using game theory to study MEV is crucial for understanding participant behavior and mitigating negative externalities in blockchain systems, as highlighted by the formalization of MEV games and the comparison of transaction ordering mechanisms in related work [29]. The integration of agents with two strategic approaches and the inherent complexities of the block-building mechanism make it arduous to directly compute the Nash equilibrium for the proposed game-theoretic model. We adopt agent-based simulation methods and genetic algorithms to model the co-evolution of agents, addressing the challenge highlighted in [30] regarding agent-based modeling in DeFi systems. Although we considered various agent learning approaches [31], we employ genetic algorithms for their strong global search capability via population diversity and stochastic mutation, enabling efficient evolution within a compact strategy set (size 20). Furthermore, our approach incorporates reinforcement learning principles by dynamically updating strategy fitness based on simulation outcomes, rather than relying on a predefined fitness function.
@@
This paper primarily focuses on the strategy selection and value allocation of profit-seeking actors involved in the block-building process. We assume that builders will use the blocks they construct to participate in an equivalent of a second-price auction, without considering any strategic behavior by builders in the block-building auction. However, a more interesting scenario is that each agent participates as a block builder in the block building auction, while they can also choose to share their bundles to reduce risk. This leads to complex bidding strategies in the block building auction. Recently, Wu et al. [24, 25] construct an agent-based model to perform EGTA on the bidding strategies of builders in the block building auction associated with the PBS system. Combining models of these two distinct stages of the MEV supply chain would induce a more complex game-theoretic problem but could offer deeper insights into understanding PBS systems.
@@ VI Conclusion
This study introduces a co-evolutionary framework to analyze the strategic behavior of profit-seeking participants in the block-building process within the PBS system. By employing genetic algorithms, we simulate the co-evolution of agents’ strategies, enabling us to observe their behavioral dynamics under varying conditions, particularly changes in the probabilities of conflicts between bundles. Through agent-based simulations, we capture the evolutionary outcomes of the system and apply EGTA, specifically the $\alpha$-Rank algorithm, to compute the selection frequencies of agents for two key meta-strategies: block building and bundle sharing.